OpenStudy (anonymous):

Simplify: (-5)(-4)(-3)(-2)(-1)

OpenStudy (anonymous):

-120

OpenStudy (anonymous):

5 negatives means a negative answer. 5*4=20.. 20*3=60.. 60*2 = 120.. *-1 = -120.